# A Review of Bayesian Uncertainty Quantification in Deep Probabilistic Image Segmentation

### v4

@@ 6.3 Model introspection
Erroneous predictions should correlate with the uncertainty in cases the source is assumed to be epistemic. In turn, this implies that the provided uncertainty can be used to to gauge prediction quality of a model. High uncertainty predictions can be discarded to ensure a level of model reliability. Hence, the described problem also be considered as Out-of-Distribution (OOD) detection (Lambert et al., 2022; Holder & Shafique, 2021). Including information on localized uncertainty to the training objective has shown to improve generalization capabilities (Ozdemir et al., 2017; Bian et al., 2020; Iwamoto et al., 2021; Li et al., 2021). Note that using uncertainty to guide model training is closely related to Active Learning, which has been discussed in Section 6.2.
@@
Considering Table 3, it can be obseverd that epistemic methods such as MC dropout and Ensemblign are by far the most popul;ar choice of uncertainty encapsulation method. However, Concrete dropout (Rumberger et al., 2020b; Mukhoti & Gal, 2018), M-heads (auxiliary networks) (Jungo & Reyes, 2019; Linmans et al., 2020), MoE (Pavlitskaya et al., 2020) Laplace Approximation (Zepf et al., 2023b), Variational Inference in 3D (LaBonte et al., 2019), and test-time augmentation (Dahal et al., 2020) have also been used for this application.
@@ 6.5 Application domains
Brain MRI Philps et al. (2024); Bhat et al. (2023); Hu et al. (2023); Philps et al. (2024); Savadikar et al. (2021); Zhang et al. (2022c) Zepf et al. (2023b); Ma et al. (2024b); Shen et al. (2021) Roy et al. (2018b); Mehrtash et al. (2020); Jungo & Reyes (2019); Eaton-Rosen et al. (2018); Wang et al. (2019b); Jungo et al. (2020); Whitbread & Jenkinson (2022); Monteiro et al. (2020); Lambert et al. (2022); Jungo et al. (2018b; a) Wolleb et al. (2021); Wu et al. (2023b; a); Chen et al. (2023); Wang et al. (2019a); Kamnitsas et al. (2018); Larrazabal et al. (2021b); Wang et al. (2019b); Whitbread & Jenkinson (2022)
@@
Lung CT Rafael-Palou et al. (2021); Gao et al. (2022); Valiuddin et al. (2024b); Chen et al. (2023); Zhang et al. (2022b); Monteiro et al. (2020); Qiu & Lui (2020); Kohl et al. (2019a); Hu et al. (2022); Zhang et al. (2022c); Gantenbein et al. (2020); Viviers et al. (2023b; c); Valiuddin et al. (2024a); Zbinden et al. (2023); Zepf et al. (2024); Kahl et al. (2024); Bhat et al. (2022a); Long et al. (2021b); Valiuddin et al. (2021); Rahman et al. (2023); Kohl et al. (2019b); Hu et al. (2023); Bhat et al. (2023); Kassapis et al. (2021); Long et al. (2021a); Selvan et al. (2020); Baumgartner et al. (2019) Kahl et al. (2024); Hoebel et al. (2020) Kahl et al. (2024); Czolbe et al. (2021); Chotzoglou & Kainz (2019) Hu et al. (2022); Rakic et al. (2024); Zhang et al. (2022a)
@@ 7 Discussion
  • Active Learning is a promising application of uncertainty quantification, reducing annotation costs while maintaining or even improving model generalization (Section 7.4)..
@@
  • Alternatives to epistemic uncertainty modeling warrant further exploration due to longstanding criticisms of the predominantly used Monte Carlo dropout and its ability to capture true model uncertainty (Section 7.7).
@@ 7.1 Disentangling uncertainties
Proper disentanglement of the uncertainties allows to accredit high-entropy outputs to either model ignorance about the data (epistemic), or inherent noise in the data generation process (aleatoric). Importantly, this distinction relates to the suggestion whether the uncertainty is reducible or not. Furthermore, Equation( 2) indicates this can be achieved by introducing stochasticity into the model parameters or the features. Seemingly straightforward, sufficient literature and discussion in the scientific community exist the applicability of this taxonomy in practical cases . As paraphrased from Kirchhof et al. (2025), definitions of uncertainty resemble the shapes of clouds — clearly defined from a distance, but losing clarity during approach by dissolving in one another. We further explore this concept and show how confusion and contradiction have led to common pitfalls in related studies. To address these issues, we highlight key nuances in uncertainty modeling and contextualize this to segmentation problems by providing recommendations for future research.
@@ 7.1.1 Nuances in uncertainty modeling
It is essential to consider that interpretation of the uncertainty source often differs depending on the modeling context. For example, epistemic uncertainty to be considered as the number of required models to fit the data (Wimmer et al., 2023), evident in ensembling and M-head approaches for tasks related to model introspection and generalization (Section 6.3 and 6.4. Another perspective pertains model disagreement (Houlsby et al., 2011; Gal et al., 2017b; Kirsch, 2024), relating to the source of observer variability (Section 6.1) and methods that employ mixture-of-expert approaches (Section 5.1.2). Also, it has been simply considered to be the remainder after subtracting the aleatoric from the total uncertainty Depeweg et al. (2018), such as commonly done in Active Learning (Section 6.2). Aleatoric uncertainty also knows several definitions such as being the Bayes-optimal residual risk Apostolakis (1990); Helton (1997), but remains consistent in related literature as the inherent, irreducible noise in the data inducing variance in the ground-truth distribution (Section 6.1). However, it was highlighted that the decision to model a specific type of uncertainty was largely influenced by how the problem was framed. To avoid confusion. this task-dependency should be noted when interpreting uncertainty.
@@
Moreover, past literature as well as our overview indicate the influence of specific design choices on the nature of the modeled uncertainty (Der Kiureghian & Ditlevsen, 2009). For instance, the expected entropy w.r.t. a stochastic parameter variable should represent the aleatoric uncertainty (Equation (3). However, when employing generative models, this behavior changes. The latent (feature) stochasticity variable is employed to model the aleatoric uncertainty. Hence, the remaining uncertainty, i.e. epistemic, is encapsulated in the expected entropy w.r.t. this latent variable (Equation (13). Give this, the calibration network of the GAN-based CAR model of Kassapis et al. (2021) (discussed in Section 4.2.1) actually quantified the epistemic rather than aleatoric uncertainty. This is dependency is also demonstrated by the conflicting ideas on uncertainty encapsulated with TTA (Wang et al., 2019a; Ayhan & Berens, 2022; Zhang et al., 2022a; Wang et al., 2019b; Rakic et al., 2024; Kahl et al., 2024; Hu et al., 2022; Whitbread & Jenkinson, 2022; Roshanzamir et al., 2023). The key message thus far is that, after selecting the type of uncertainty to model, it is equally important to carefully choose the appropriate method for its representation, as confusion and misunderstandings can cause deviation from the targeted uncertainty type.
@@
Despite the seemingly rigorous and well-defined distinction, it can be seen that implementation rapidly blurs this notion. A systematic study has found that the ability to separate the uncertainties can strongly depend on the data Kahl et al. (2024). These findings are also supported by research presenting a strong correlation between epistemic and aleatoric uncertainty quantification methods de Jong et al. (2024); Mucsányi et al. (2024), indicating that they represent uncertainty of similar kind. Kendall & Gal (2017) have also observed that explicitly modeling a single uncertainty tends to compensate for the lack of the other. In probabilistic models, we can also find that the epistemic encapsulation methods have been used to model observer variability. Furthermore, using a combination of aleatoric end epistemic modeling techniques can also beneficial (Gao et al., 2022). These finding further highlight the notion of the uncertainties being intertwined. On a slight side note, it is notable that popular BNN methods such as the Local Reparameterization trick use stochasticity in the output features of intermediate layers (usually considered to encapsulate aleatoric uncertainty) as a proxy to model the weight distributions Kingma et al. (2015), further supporting this argument.
@@
A common misconception pertains that modeling a specific uncertainty also implies that the predicted output represents that same uncertainty. This undermines the entropy decomposition in Equation (3) distinguishing between encapsulating, i.e. targeting an intended uncertainty type, versus quantifying, i.e. measuring the modeled uncertainty in isolation. For example, the estimation of mutual information is required to express the epistemic uncertainty in isolation. Nonetheless it is common for papers to quantify the uncertainty by inspecting the output distribution subject through repetitive sampling from the parameters densities, thereby approximating the predictive entropy encapsulating both uncertainties(Camarasa et al., 2021; Seeböck et al., 2019; Roshanzamir et al., 2023; Roy et al., 2018a; Rumberger et al., 2020b; Hoebel et al., 2022; Seeböck et al., 2019; Wundram et al., 2024b; Rumberger et al., 2020a; Sedai et al., 2018). At the very most, the output correlates to an increase in epistemic uncertainty. However, given the entanglement of both uncertainties, part of the predictive variance can be accredited to noise in the data. This similarly holds for literature emphasizing aleatoric uncertainty when estimating the predictive variance Zhang et al. (2022c); Monteiro et al. (2020); Valiuddin et al. (2024b); Gao et al. (2022). However, epistemic uncertainty is often assumed to be explained away or assumed to be comparatively negligible with large datasets. Finally, we have found it to be common to interchangeably use the total and model uncertainty, especially in the context of Monte Carlo Dropout methods Kendall et al. (2016); Roy et al. (2018b); DeVries & Taylor (2018); Mehrtash et al. (2020); Burmeister et al. (2022); Iwamoto et al. (2021). To prevent confusion in already established terminology, the term model uncertainty should remain reserved for the induced ambiguity due to limited model complexity, rather than to denote general or the entire epistemic uncertainty.
@@ 7.1.2 Pragmatic use of uncertainty
As argued by Kahl et al. (2024), claiming to quantify and uncertainty type would ideally require validation across all relevant downstream tasks. For example, validation of epistemic uncertainty requires a clear distribution shift in the data (domain, semantic, covariate etc.), but is found in literature that often either no distribution shift or the required metrics are used (Zhang et al., 2022a; Wang et al., 2019a; Postels et al., 2019; Mukhoti & Gal, 2018; Mobiny et al., 2021; Whitbread & Jenkinson, 2022). As discussed in Section 6.1, aleatoric uncertainty should be validated with multiple labels and appropriate metrics, but can often found to be evaluated on a single annotator or without the use of distribution-level statistics Wang et al. (2019a); Whitbread & Jenkinson (2022); Kendall & Gal (2017); Liu et al. (2022b); Schmidt et al. (2023); Savadikar et al. (2021). Hence, simply avoiding any serious claims related to either type alleviates this burden, and allows one to focus solely on the task at hand. In contrast to the work of Kahl et al. (2024), however, we do not consider calibration of models as a downstream task, but rather as a tool to quantify model reliability. As a consequence, the pixel-wise uncertainties can be used for possible downstream applications.
@@ 7.2 Spatial coherence
Quantifying the epistemic uncertainty requires subtracting the expected entropy from the total entropy of the predictive distribution (see equations (3) and (39)). In such cases, the entropy is usually estimated via mean or summation over all pixel-level uncertainty values Mukhoti & Gal (2018); Camarasa et al. (2021); Bhat et al. (2022b); Mukhoti et al. (2020); Ma et al. (2024b); Shen et al. (2021); Li & Alstrøm (2020); Dechesne et al. (2021); Huang et al. (2018); Nair et al. (2020); Zepf et al. (2023b). However, this asserts the assumption of a factorized categorical distribution, i.e. uncorrelated pixels, which is a strong assumption that ignores the core rationale for the field’s existence. Since this could explain failure cases of entropy-based uncertainty evaluation methods, it is important to shift towards contemporary generative models that enable exact likelihoods, improving the theoretical soundness of the evaluated entropies. Moreover, elements such as object class dependency, size, shape, proximity, contextual information etc. can all influence the decision-making of the aggregation method. Notably, Kahl et al. (2024) argue that pixel aggregation should be considered as a separate, distinct modeling choice in probabilistic segmentation. The authors demonstrate that naive summation-based aggregation techniques (Wang et al., 2019a; Jungo et al., 2020; Zhang et al., 2022a; Czolbe et al., 2021; Camarasa et al., 2021) can cause the foreground object size to correlate with the uncertainty score. Fortunately, some empirical studies have explored aggregation (Camarasa et al., 2021; Roy et al., 2018b; Jungo et al., 2020; Kasarla et al., 2019; Xie et al., 2022; Wu et al., 2021). However, we advocate for deeper investigation into well-informed image-level aggregation strategies and their implications for uncertainty quantification.
@@
In clinical settings, data is often acquired volumetrically but sliced into 2D patches due to memory limits. With improved compute, 3D segmentation models have gained traction for preserving spatial continuity and reducing inter-slice inconsistencies. As we have seen the implications of the additional dimension in 2D uncertainty, we similarly argue 3D segmentation to introduce or exacerbate previously mentioned issues related to entropy estimation and aggregation of pixels. Probabilistic segmentation has been employed on 3D data but often are straightforward extensions of patch-wise segmentation architectures (Viviers et al., 2023b; Chotzoglou & Kainz, 2019; Long et al., 2021b; Saha et al., 2021a; 2020; Hoebel et al., 2019). Limited diversity of uncertainty-based sample selection in Active Learning requires more focus, since 2D slices of 3D patches are often highly redundant Shen et al. (2021); Burmeister et al. (2022). Furthermore, given the high computational cost of many architectures, we emphasize the need for efficiency with higher-dimensional input tensors. Furthermore, methodologies need to be developed in order to appropriately compare 2D to 3D models. For example, the GED cannot be estimated volumetrically with 2D models. Hence, general guidelines related to translating 2D segmentation models to 3D can be of great benefit for practitioners considering to phase out patch-based approaches for volumetric data. Fortunately, research dedicated to probabilstic volumetric segmentation is prevalent and its successes underline the potential for this research direction Liu et al. (2025); Burmeister et al. (2022); Viviers et al. (2023b).

### v5

@@ 6.3 Model introspection
Erroneous predictions should correlate with the uncertainty in cases the source is assumed to be epistemic. In turn, this implies that the provided uncertainty can be used to to gauge prediction quality of a model. High uncertainty predictions can be discarded to ensure a level of model reliability. Hence, the described problem can also be considered as Out-of-Distribution (OOD) detection (Lambert et al., 2022; Holder & Shafique, 2021). Including information on localized uncertainty to the training objective has shown to improve generalization capabilities (Ozdemir et al., 2017; Bian et al., 2020; Iwamoto et al., 2021; Li et al., 2021). Note that using uncertainty to guide model training is closely related to Active Learning, which has been discussed in Section 6.2.
@@
Considering Table 3, it can be observed that epistemic methods such as MC dropout and Ensembles are by far the most popular choice of uncertainty encapsulation method. However, Concrete dropout (Rumberger et al., 2020b; Mukhoti & Gal, 2018), M-heads (auxiliary networks) (Jungo & Reyes, 2019; Linmans et al., 2020), MoE (Pavlitskaya et al., 2020) Laplace Approximation (Zepf et al., 2023b), Variational Inference in 3D (LaBonte et al., 2019), and test-time augmentation (Dahal et al., 2020) have also been used for this application.
@@ 6.5 Application domains
Brain MRI Philps et al. (2024); Bhat et al. (2023); Hu et al. (2023); Philps et al. (2024); Savadikar et al. (2021); Zhang et al. (2022c) Zepf et al. (2023b); Ma et al. (2024b); Shen et al. (2021) Roy et al. (2018b); Mehrtash et al. (2020); Jungo & Reyes (2019); Eaton-Rosen et al. (2018); Wang et al. (2019b); Jungo et al. (2020); Whitbread & Jenkinson (2022); Monteiro et al. (2020); Lambert et al. (2022); Jungo et al. (2018b; a) Wolleb et al. (2021); Wu et al. (2023b; a); Chen et al. (2023a); Wang et al. (2019a); Kamnitsas et al. (2018); Larrazabal et al. (2021b); Wang et al. (2019b); Whitbread & Jenkinson (2022)
@@
Lung CT Rafael-Palou et al. (2021); Gao et al. (2022); Valiuddin et al. (2024b); Chen et al. (2023a); Zhang et al. (2022b); Monteiro et al. (2020); Qiu & Lui (2020); Kohl et al. (2019a); Hu et al. (2022); Zhang et al. (2022c); Gantenbein et al. (2020); Viviers et al. (2023b; c); Valiuddin et al. (2024a); Zbinden et al. (2023); Zepf et al. (2024); Kahl et al. (2024); Bhat et al. (2022a); Long et al. (2021b); Valiuddin et al. (2021); Rahman et al. (2023); Kohl et al. (2019b); Hu et al. (2023); Bhat et al. (2023); Kassapis et al. (2021); Long et al. (2021a); Selvan et al. (2020); Baumgartner et al. (2019) Kahl et al. (2024); Hoebel et al. (2020) Kahl et al. (2024); Czolbe et al. (2021); Chotzoglou & Kainz (2019) Hu et al. (2022); Rakic et al. (2024); Zhang et al. (2022a)
@@ 7 Discussion
  • Active Learning is a promising application of uncertainty quantification, reducing annotation costs while maintaining or even improving model generalization (Section 7.4).
@@
  • Alternatives to epistemic uncertainty modeling warrant further exploration due to longstanding criticisms of the predominantly used Monte Carlo dropout, questioning its ability to capture true model uncertainty (Section 7.7).
@@ 7.1 Disentangling uncertainties
Proper disentanglement of the uncertainties allows to accredit high-entropy outputs to either model ignorance about the data (epistemic), or inherent noise in the data generation process (aleatoric). Importantly, this distinction relates to the suggestion whether the uncertainty is reducible or not. Furthermore, Equation( 2) indicates this can be achieved by introducing stochasticity into the model parameters or the features. Seemingly straightforward, sufficient literature and discussion in the scientific community exist related to the applicability of this taxonomy in practical cases. As paraphrased from Kirchhof et al. (2025), “definitions of uncertainty resemble the shapes of clouds — clearly defined from a distance, but losing clarity during approach by dissolving in one another”. We further explore this concept and show how confusion and contradiction have led to common pitfalls in related studies. To address these issues, we highlight key nuances in uncertainty modeling and contextualize this to segmentation problems by providing recommendations for future research.
@@ 7.1.1 Nuances in uncertainty modeling
It is essential to consider that interpretation of the uncertainty source often differs depending on the modeling context. For example, epistemic uncertainty to be considered as the number of required models to fit the data (Wimmer et al., 2023), evident in ensembling and M-head approaches for tasks related to model introspection and generalization (Sections 6.3 and 6.4). Another perspective pertains model disagreement (Houlsby et al., 2011; Gal et al., 2017b; Kirsch, 2024), relating to the source of observer variability (Section 6.1) and methods that employ mixture-of-expert approaches (Section 5.1.2). Also, it has been simply considered to be the remainder after subtracting the aleatoric from the total uncertainty Depeweg et al. (2018), such as commonly done in Active Learning (Section 6.2). Aleatoric uncertainty also knows several definitions such as being the Bayes-optimal residual risk Apostolakis (1990); Helton (1997), but remains consistent in related literature as the inherent, irreducible noise in the data inducing variance in the ground-truth distribution (Section 6.1). However, it was highlighted that the decision to model a specific type of uncertainty was largely influenced by how the problem was framed. To avoid confusion, this task-dependency should be noted when interpreting uncertainty.
@@
Moreover, past literature as well as our overview indicate the influence of specific design choices on the nature of the modeled uncertainty (Der Kiureghian & Ditlevsen, 2009). For instance, the expected entropy w.r.t. a stochastic parameter variable should represent the aleatoric uncertainty (Equation (3)). However, when employing generative models, this behavior changes. The latent (feature) stochasticity variable is employed to model the aleatoric uncertainty. Hence, the remaining uncertainty, i.e. epistemic, is encapsulated in the expected entropy w.r.t. this latent variable (Equation (13)). Hence, the calibration network of the GAN-based CAR model of Kassapis et al. (2021) (discussed in Section 4.2.1) actually quantified the epistemic rather than aleatoric uncertainty. This is dependency is also demonstrated by the conflicting ideas on uncertainty encapsulated with TTA (Wang et al., 2019a; Ayhan & Berens, 2022; Zhang et al., 2022a; Wang et al., 2019b; Rakic et al., 2024; Kahl et al., 2024; Hu et al., 2022; Whitbread & Jenkinson, 2022; Roshanzamir et al., 2023). The key message thus far is that, after selecting the type of uncertainty to model, it is equally important to carefully choose the appropriate method for its representation, as confusion and misunderstandings can cause deviation from the targeted uncertainty type.
@@
Despite the seemingly rigorous and well-defined distinction, it can be seen that implementation rapidly blurs this notion. A systematic study has found that the ability to separate the uncertainties can strongly depend on the data Kahl et al. (2024). These findings are also supported by research presenting a strong correlation between epistemic and aleatoric uncertainty quantification methods de Jong et al. (2024); Mucsányi et al. (2024), indicating that they represent uncertainty of similar kind. Kendall & Gal (2017) have also observed that explicitly modeling a single uncertainty tends to compensate for the lack of the other. In probabilistic models, we can also find that the epistemic encapsulation methods have been used to model observer variability. Furthermore, using a combination of aleatoric end epistemic modeling techniques can also beneficial (Gao et al., 2022). These findings further highlight the notion of the uncertainties being intertwined. On a slight side note, it is notable that popular BNN methods such as the Local Reparameterization trick use stochasticity in the output features of intermediate layers (usually considered to encapsulate aleatoric uncertainty) as a proxy to model the weight distributions Kingma et al. (2015), further supporting this argument.
@@
A common misconception pertains that modeling a specific uncertainty also implies that the predicted output represents that same uncertainty. This undermines the entropy decomposition in Equation (3) distinguishing between encapsulating, i.e. targeting an intended uncertainty type, versus quantifying, i.e. measuring the modeled uncertainty in isolation. For example, the estimation of mutual information is required to express the epistemic uncertainty in isolation. Nonetheless it is common for papers to quantify the uncertainty by inspecting the output distribution subject through repetitive sampling from the parameters densities, thereby approximating the predictive entropy encapsulating both uncertainties (Camarasa et al., 2021; Seeböck et al., 2019; Roshanzamir et al., 2023; Roy et al., 2018a; Rumberger et al., 2020b; Hoebel et al., 2022; Seeböck et al., 2019; Wundram et al., 2024b; Rumberger et al., 2020a; Sedai et al., 2018). At the very most, the output correlates with an increase in epistemic uncertainty. However, given the entanglement of both uncertainties, part of the predictive variance can be accredited to noise in the data. This similarly holds for literature emphasizing aleatoric uncertainty when estimating the predictive variance Zhang et al. (2022c); Monteiro et al. (2020); Valiuddin et al. (2024b); Gao et al. (2022). However, epistemic uncertainty is often assumed to be explained away or assumed to be comparatively negligible with large datasets. Finally, we have found it to be common to interchangeably use the total and model uncertainty, especially in the context of Monte Carlo Dropout methods Kendall et al. (2016); Roy et al. (2018b); DeVries & Taylor (2018); Mehrtash et al. (2020); Burmeister et al. (2022); Iwamoto et al. (2021). To prevent confusion in already established terminology, the term model uncertainty should be exclusively used for the induced ambiguity due to limited model complexity, rather than to denote general or the entire epistemic uncertainty.
@@ 7.1.2 Pragmatic use of uncertainty
As argued by Kahl et al. (2024), claiming to quantify an uncertainty type would ideally require validation across all relevant applications. For example, validation of epistemic uncertainty requires a clear distribution shift in the data (domain, semantic, covariate etc.), but it is often found in literature that either no distribution shift or the required metrics are used (Zhang et al., 2022a; Wang et al., 2019a; Postels et al., 2019; Mukhoti & Gal, 2018; Mobiny et al., 2021; Whitbread & Jenkinson, 2022). As discussed in Section 6.1, aleatoric uncertainty should be validated with multiple labels and appropriate metrics, but can often found to be evaluated on a single annotator or without the use of distribution-level statistics Wang et al. (2019a); Whitbread & Jenkinson (2022); Kendall & Gal (2017); Liu et al. (2022b); Schmidt et al. (2023); Savadikar et al. (2021). Hence, simply avoiding any serious claims related to either type alleviates this burden, and allows one to focus solely on the task at hand. In contrast to the work of Kahl et al. (2024), however, we do not consider calibration of models itself as a downstream task, but rather as a tool to gauge model reliability to be later used for other tasks. As a consequence, the pixel-wise uncertainties can be later used for possible downstream applications.
@@ 7.2 Spatial coherence
Quantifying the epistemic uncertainty requires subtracting the expected entropy from the total entropy of the predictive distribution (see equations (3) and (39)). In such cases, the entropy is usually estimated via mean or summation over all pixel-level uncertainty values Mukhoti & Gal (2018); Camarasa et al. (2021); Bhat et al. (2022b); Mukhoti et al. (2020); Ma et al. (2024b); Shen et al. (2021); Li & Alstrøm (2020); Dechesne et al. (2021); Huang et al. (2018); Nair et al. (2020); Zepf et al. (2023b). However, this asserts the assumption of a factorized categorical distribution, i.e. uncorrelated pixels, which is a strong assumption that ignores the core rationale for the field’s existence. Since this could explain failure cases of entropy-based uncertainty evaluation methods, it is important to shift towards contemporary generative models that enable exact likelihoods, improving the theoretical soundness of the evaluated entropies. Moreover, elements such as object class dependency, size, shape, proximity, contextual information etc. can all influence the decision-making of the aggregation method. Notably, Kahl et al. (2024) argue that pixel aggregation should be considered as a separate, distinct modeling choice in probabilistic segmentation. The authors demonstrate that naive summation-based aggregation techniques (Wang et al., 2019a; Jungo et al., 2020; Zhang et al., 2022a; Czolbe et al., 2021; Camarasa et al., 2021) can cause the foreground object size to correlate with the uncertainty score. Some empirical studies have explored aggregation (Camarasa et al., 2021; Roy et al., 2018b; Jungo et al., 2020; Kasarla et al., 2019; Xie et al., 2022; Wu et al., 2021). However, we advocate for deeper investigation into well-informed image-level aggregation strategies and their implications for uncertainty quantification.
@@
In clinical settings, data is often acquired volumetrically but sliced into patches due to memory limits. With improved compute, 3D segmentation models have gained traction for preserving spatial continuity and reducing inter-slice inconsistencies. As we have seen the implications of the additional dimension in 2D uncertainty, we similarly argue that 3D segmentation introduces or exacerbates previously mentioned issues related to entropy estimation and aggregation of pixels. Probabilistic segmentation has been employed on 3D data but often are straightforward extensions of patch-wise segmentation architectures (Viviers et al., 2023b; Chotzoglou & Kainz, 2019; Long et al., 2021b; Saha et al., 2021a; 2020; Hoebel et al., 2019). Limited diversity of uncertainty-based sample selection in Active Learning requires more focus, since 2D slices of 3D volumes are often highly redundant Shen et al. (2021); Burmeister et al. (2022). Furthermore, given the high computational cost of many architectures, we emphasize the need for efficiency with higher-dimensional input tensors. Furthermore, methodologies need to be developed in order to appropriately compare 2D- to 3D models. For example, the GED cannot be estimated volumetrically with 2D models. Hence, general guidelines related to translating 2D segmentation models to 3D can be of great benefit for practitioners considering to phase out patch-based approaches for volumetric data. Fortunately, research dedicated to probabilistic volumetric segmentation is prevalent and its successes underline the potential for this research direction Liu et al. (2025); Burmeister et al. (2022); Viviers et al. (2023b).